A motorboat can maintain a constant speed of 16 mph relative to the water. The boat makes a trip upstream to a marina in 20 minutes. The return trip takes 15 minutes. What is the speed of the current?
Answer by mananth(16946)

Let current speed = x mph
upstream speed = 16-x mph
downstream speed= 16+x mph
Time upstream = 1/3 hours
Time downstream = 1/4 hours
Distance upstream = distance downstream
distance = speed * time
1/3(16-x) = 1/4(16+x)
multiply equation by LCD 12
4(16-x)=3(16+x)
64-4x = 48+3x
64-48= 7x
16=7x
x= 16/7
x= 2.3 mph the current speed